A firm is analysing the impact of a price increase on frozen vegetables. The price was increased from R35 to R38 per kilogramme, resulting in a decline in quantity demanded, from 1 400 to 1 350 kilogrammes per day.

2.1 What is the price elasticity of demand for frozen vegetables in this price range, based on the midpoint formula? (4)

2.2 Given that the price increase resulted in reduced sales, using the elasticity coefficient and the revenue test, explain whether you think the decision to increase the price was sound
1
Expert's answer
2019-08-23T15:39:51-0400

2.1 Price elasticity indicates how quantity demanded has changed due to a change in prices. In the use of midpoint formula, the price elasticity is the same regardless of the direction. This is because midpoint formula focus on the use of average value acquired from dividing initial and final value by 2.

Therefore, price elasticity using midpoint formula is given as: 

"Price Elasticity = \frac {(Q2-Q1)/[(Q2+Q1)/2]}{(P2-P1)/[(P2+P1)/2]}"


Where: Q is the quantity demanded

 P is the price

Q1=1400                                                       P1=35

Q2=1350                                                       P2=38


"Price Elasticity = \frac{(-50/1375)}{(3/36.5)}"


Price Elasticity = -0.036/0.082

                         = -0.44


2.2. From the price elasticity coefficient of -0.44, it means that from every percentage increase in price of frozen vegetables, the quantity demand decrease by 0.44. This indicates that increase in price has an insignificant impact on the quantity demanded but this impact can be proved with the use of revenue test.

In the use of revenue test;

"New Revenues (R2) =\ P2Q2"

where P2 is the new price and Q2 is the new quantity demanded

   

                                         = (38 x 1350)

                                         = R51300


"Previous Revenues (R1) = \ P1Q1"

                                               = (35x1400)

                                               = R49000


Increase in revenue = (R2-R1)

                                          = R (51300-49000)

                                          = R2300


An increase in revenue justifies that the decision to increase prices was sound.